# Human4DiT: Free-view Human Video Generation with 4D Diffusion Transformer

### v1

@@ Abstract
We present a novel approach for generating high-quality, spatio-temporally coherent human videos from a single image under arbitrary viewpoints. Our framework combines the strengths of U-Nets for accurate condition injection and diffusion transformers for capturing global correlations across viewpoints and time. The core is a cascaded 4D transformer architecture that factorizes attention across views, time, and spatial dimensions, enabling efficient modeling of the 4D space. Precise conditioning is achieved by injecting human identity, camera parameters, and temporal signals into the respective transformers. To train this model, we curate a multi-dimensional dataset spanning images, videos, multi-view data and 3D/4D scans, along with a multi-dimensional training strategy. Our approach overcomes the limitations of previous methods based on GAN or UNet-based diffusion models, which struggle with complex motions and viewpoint changes. Through extensive experiments, we demonstrate our method’s ability to synthesize realistic, coherent and free-view human videos, paving the way for advanced multimedia applications in areas such as virtual reality and animation. Our project website is https://human4dit.github.io.
@@
Figure 1: We propose Human4DiT, a novel approach to generate high-quality, spatio-temporally coherent human videos given a reference image. With the proposed 4D diffusion transformer, our method is capable of generating monocular video, multi-view video, 3D static video, and free-view video.
footnotetext: Equal Contribution.
1 Introduction
@@
Human video generation is an active topic in the field of video generation. It has widespread applications in areas such as virtual reality, animation, gaming, and movie production. Moreover, generating realistic human videos holds great significance for advancing multimedia technologies and enabling new forms of human-computer interaction.
@@
Recently, with the rapid development of diffusion models, especially latent diffusion models [24], leveraging diffusion models for human video generation has become the mainstream approach [37, 41, 50, 7]. To incorporate human priors as control conditions into the diffusion model, some approaches [11] have adopted skeleton-based schemes, using the skeletal connectivity graph as a control condition either through ControlNet [47] or by concatenating it with the input. Other methods are based on the SMPL body model template [16], injecting SMPL-derived representations such as UV maps, depth, normals [50], or direct dense pose embeddings [41] into the diffusion model. Current human video diffusion models based on the UNet architecture [25] from Stable Diffusion could inject control conditions into the network in a pixel-aligned manner. However, the UNet’s reliance on local convolutional operations makes it more focus on local generation, resulting in relatively poorer performance for the global aspects, especially when generating long and complex human motions. Moreover, these methods only consider the human body itself, neglecting viewpoint information from the camera perspective, especially for scenarios involving significant viewpoint changes such as 360-degree free-viewpoint human video generation. Incorporating viewpoint control signals into the network while simultaneously maintaining coherence across different viewpoints and time poses a significant challenge.
To overcome the challenges of generating complex human motions across views and time, we propose a novel video generation network architecture that combines the strengths of UNets and diffusion transformers. The UNet component is responsible for accurate pixel-aligned condition injection, which has been proven powerful for human image generation. However, it remains a challenge to ensure temporal consistency when using UNet-alike architectures to generate long videos. Recently, OpenAI’s recent work on long video generation, SORA [20], adopted a diffusion transformer [21] architecture and demonstrated substantially better realism and spatio-temporal coherence than UNet-based competitors. Inspried by SORA, we introduce a 4D diffusion transformer, which learns global correlations of complex human motions across views and time, thereby preserving spatio-temporal consistency of generated human videos. In practice, we first apply UNet to extract pixel-aligned features, after which these features are tokenized and processed by the 4D diffusion transformer with self-attention mechanisms.
@@
However, directly applying a diffusion transformer to learn correlations over the views and time would be computationally prohibitive. Therefore, we propose an efficient novel 4D transformer architecture. Its core principle is to cascadingly learn the correlations across the 4D space (view, time, height, width) via self-attention. Specifically, we factorize the 4D diffusion transformer into three transformer blocks: 2D image transformer blocks, temporal transformer blocks, and view transformer blocks, each attending to different dimensions of the 4D space. These three types of blocks are interconnected to form a 4D transformer block. Multiple such 4D transformer blocks are then cascaded to construct the final 4D transformer. This efficiently captures the interrelations between body parts (height, width) across viewpoints (views) and time steps (times).
@@
Based on our proposed network architecture, we incorporate different control signals into the respective network modules for precise control, including SMPL, human identity, time, and camera parameters. SMPL features and human identity embeddings are extracted by the guidance UNet and injected into all three transformers; view embeddings are encoded from camera parameters and injected into the view transformer; temporal embeddings are injected into the temporal transformer. Through these modules, we effectively inject various control conditions into the network, enabling arbitrary viewpoint control and the generation of high-quality, coherent human videos.
@@
To train our proposed 4D diffusion transformer model, we also collected a large multi-dimensional dataset and devised a multi-dimensional training strategy that fully leverages all available data modalities. Our multi-dimensional dataset comprises images, videos, multi-view videos, 3D scans, as well as a limited amount of 4D scans spanning different viewpoints and time steps. During the inference stage, we propose a spatio-temporally consistent diffusion sampling strategy that enables the generation of long free-viewpoint 360-degree videos despite limited spatio-temporal window constraints. The strategy is carried out with two stages. The first stage treats the free-view video as a monocular long video sequence, maximizing the temporal window to ensure long-term temporal consistency. After that, the second stage regards the free-view video as a collection of multi-view video clips, using a larger viewpoint window with a smaller temporal window, encouraging consistency across viewpoints.
To summarize, our main contributions are threefold:
@@
  • We are the first to introduce diffusion transformers to human video generation. Combined with methods that simply use UNet and several control modules, our method achieves higher-quality free-viewpoint and spatio-temporally consistent generation of long human videos.
@@
  • During inference, we also propose a spatio-temporally consistent diffusion sampling strategy to generate coherent 360-degree long human videos.
2 Related Work
2.1 Human Video Generation
@@
GAN-based methods [30, 32, 38, 39, 28, 29] leverage the inherent generative capabilities of adversarial networks [5, 19] to spatially transform reference images according to input motion. These approaches commonly employ warping functions to generate sequential video frames, aiming to fill in missing regions and improve visually implausible areas within the generated content. While GAN-based methods have shown promising results in dynamic visual content generation, they often struggle with effectively transferring motion, especially when there are significant variations in human identity and scene dynamics between the reference image and the source video motion. This can lead to unrealistic visual artifacts and temporal inconsistencies in the synthesized videos.
@@
On the other hand, diffusion models, known for the superior generation quality and stable controllability, have been successfully integrated into human image animation. These models [2, 10, 37, 7, 41, 50] employ various strategies, such as texture diffusion blocks, optical flow synthesis in latent space, and motion representation using flow maps, to enhance the visual fidelity of the generated videos. Animate Anyone [7] employs a UNet-based ReferenceNet to extract features from reference images and incorporates pose information through an efficient pose guider. However, recent diffusion-based methods are still mainly relying on image diffusion model and face challenges in maintaining texture consistency and temporal stability across frames. Furthermore, these methods don’t explore view controllability.
2.2 Diffusion Model with Camera Control
@@
Camera information is conditioned in diffusion models for view control. One line of works inject camera parameters into the text-to-image diffusion models for consistent view synthesis from a single input image, which has inspired researchers to explore their potential for generating consistent multi-view images. Zero-1-to-3 [13] finetunes Stable diffusion by conditioning camera poses for zero-shot novel view synthesis. Syncdreamer [14] employs 3D volumes and depth-wise attention to maintain consistency across views. MVDream [27] and other methods like Wonder3D [15] and Zero123++ [26] leverage 3D self-attention to extend multi-view image generation to more general and efficient. There are other works [42, 6] inject camera information into text-to-video (T2V) models. For example, Direct-a-Video [42] injects quantitative camera movements into temporal cross-attention blocks for view control.
2.3 Diffusion Transformer
@@
The transformer architecture [36] has revolutionized the field of natural language processing, with models like GPT [22, 23] achieving remarkable success. Recent research has demonstrated the potential of transformers in various computer vision tasks, including image classification [33, 45], semantic segmentation [49, 40, 31].
@@
Building upon this progress, the Diffusion Transformer (DiT) [21] and its variants [1, 48] have taken a step further by substituting the conventional convolutional-based U-Net backbone [25] with transformers in diffusion models. This architectural shift offers enhanced scalability compared to U-Net-based models, enabling the seamless expansion of model parameters. Recently, the transformer architecture is also integrated into text-to-video models [17, 20, 18], improving generation performance.
Figure 2: Pipeline of Human4DiT: our framework is based on 4D diffusion transformer, which adopts a cascaded structure consisting of the 2D image, the view transformer, and the temporal blocks. The input contains a reference image $\mathbf{y}_{r}$, dynamic SMPL sequences $\mathbf{P}$, and camera parameters $\mathbf{c}$. Starting from a generated noisy latent representation $x_{t}$, we then denoise $x_{t}$ conditioned on $\mathbf{y}_{r}$, $\mathbf{P}$, and $\mathbf{c}$. Firstly, the 2D image transformer block is designed to capture spatial self-attention within each frame in $x_{t}$ and $\mathbf{P}$. In addition, $y_{id}$ extracted from $\mathbf{y}_{r}$ is also injected to ensure identity consistency. Secondly, we use the view transformer block to learn correspondences across different viewpoints conditioned on $\mathbf{c}$. Finally, we adopt a temporal transformer to capture temporal correlations with time embedding. The human image embedding $y_{e}$ extracted through CLIP is shared by three blocks.
3 Overview
@@
Given a reference image of a person $\mathbf{y}_{r}$, a sequence of dynamic SMPL models $\{\mathbf{P}_{i=1}^{T}=\mathbf{(\theta,\beta)}_{i=1}^{T}\}$, and camera parameters $\{\mathbf{c}_{i=1}^{V}\}$, the goal of our method is to generate a video of that person performing the corresponding motion from the specified viewpoint. The overall pipeline of our approach is illustrated in Figure 2. Since our framework is based on a latent diffusion transformer, we first generate a noisy latent representation $\mathbf{x}_{t}$. We then inject $\mathbf{y}_{r}$, $\{\mathbf{P}_{i=1}^{T}\}$, and $\{\mathbf{c}_{i=1}^{V}\}$ as control signals into the diffusion transformer, which iteratively performs denoising to ultimately generate the latent video $\mathbf{x}_{0}$ (Sec. 4).
@@
During the inference stage, to enable the generation of free-viewpoint long-duration human motion videos under limited temporal and viewpoint window constraints, we propose an efficient diffusion sampling strategy. This strategy achieves improved spatio-temporal coherence by planning viewpoint and temporal window sizes of different scales in the diffusion sampling process (Sec. 6).
4 Network Structure
@@
In this section, we introduce the core components of our network structure: the 4D diffusion transformer and the control condition injection modules. To efficiently establish spatio-temporal relationships, our 4D diffusion transformer adopts a cascaded structure consisting of the 2D image, the temporal, and the view transformer blocks. Additionally, our control condition injection modules comprise the reference image injection module, the SMPL injection module, the time injection module, and the viewpoint injection module. Notably, the reference image injection module and the SMPL injection module utilize UNet architectures to ensure pixel-aligned injection of the control conditions.
4.1 4D Diffusion Transformer
Our 4D diffusion transformer performs denoising on a randomly initialized noisy multi-view latent human video $\mathbf{x}_{t}$. After multiple denoising steps, it produces the denoised output $\mathbf{x}_{0}$, which is eventually decoded by a VAE to generate the video. The input $\mathbf{x}_{t}$ could be regarded as a set of tokens in 5D ($\mathbf{x}_{t}\in\mathbb{R}^{V\times T\times H\times W\times C}$), where $C$ represents the feature dimension of each token, $V$ is the viewpoint dimension, $T$ is the time dimension, and $H$ and $W$ are the height and width, respectively. Our 4D diffusion transformer incorporates three types of transformer blocks to efficiently establish correlations across viewpoints, time, and 2D spatial dimensions.
First, we feed $x_{t}$ into a 2D image transformer block to capture spatial self-attention within each frame:
$\begin{split}\mathbf{x}_{t}^{s}&=\text{rearrange}(\mathbf{x}_{t},(V\times T,H\times W,C)),\\ \mathbf{Q}_{s},\mathbf{K}_{s},\mathbf{V}_{s}&=f_{s}^{\mathbf{Q}}(\mathbf{x}_{t}^{s}),f_{s}^{\mathbf{K}}(\mathbf{x}_{t}^{s}),f_{s}^{\mathbf{V}}(\mathbf{x}_{t}^{s}),\\ \mathbf{\hat{x}}_{t}^{s}&=\text{softmax}(\frac{\mathbf{Q}_{s}\mathbf{K}_{s}^{T}}{\sqrt{d_{k}}})\mathbf{V}_{s},\end{split}$ (1)
where $f_{s}^{\mathbf{Q}},f_{s}^{\mathbf{K}},f_{s}^{\mathbf{V}}$ are the linear layers in the transformer block. The rearrange operation reshapes the tensor $\mathbf{x}_{t}$ into a 3D tensor $\mathbf{x}_{t}^{s}\in R^{(V\times T)\times(H\times W)\times C}$, where the first dimension $V\times T$ is treated as the batch size. The actual attention operation is then performed on the last two dimensions $H\times W$ and $C$, which correspond to the 2D spatial dimensions of the images. After 2D image transformer, we further model correspondences across different viewpoints through the view transformer block:
$\begin{split}\mathbf{x}_{t}^{v}&=\text{rearrange}(\mathbf{\hat{x}}_{t}^{s},(T,V\times H\times W,C)),\\ \mathbf{Q}_{v},\mathbf{K}_{v},\mathbf{V}_{v}&=f_{v}^{\mathbf{Q}}(\mathbf{x}_{t}^{v}),f_{v}^{\mathbf{K}}(\mathbf{x}_{t}^{v}),f_{v}^{\mathbf{V}}(\mathbf{x}_{t}^{v}),\\ \mathbf{\hat{x}}_{t}^{v}&=\text{softmax}(\frac{\mathbf{Q}_{v}\mathbf{K}_{v}^{T}}{\sqrt{d_{k}}})\mathbf{V}_{v},\end{split}$ (2)
In the view transformer, due to the substantial variations across different viewpoints, especially between frontal, side, and back views, we perform attention jointly across the viewpoint and the 2D spatial dimensions. This attention allows for better modeling of the global correlations across the entire frame. It also makes the view transformer the most computationally and memory intensive component, limiting the maximum viewpoint window size. We will address this constraint during the inference process, as discussed later. After the view transformer, we finally employ a temporal transformer to capture temporal correlations across time steps.
$\begin{split}\mathbf{x}_{t}^{m}&=\text{rearrange}(\mathbf{\hat{x}}_{t}^{v},(V\times H\times W,T,C)),\\ \mathbf{Q}_{m},\mathbf{K}_{m},\mathbf{V}_{m}&=f_{m}^{\mathbf{Q}}(\mathbf{x}_{t}^{m}),f_{m}^{\mathbf{K}}(\mathbf{x}_{t}^{m}),f_{m}^{\mathbf{V}}(\mathbf{x}_{t}^{m}),\\ \mathbf{\hat{x}}_{t}^{m}&=\text{softmax}(\frac{\mathbf{Q}_{m}\mathbf{K}_{m}^{T}}{\sqrt{d_{k}}})\mathbf{V}_{m},\end{split}$ (3)
The three transformer blocks (2D image, view, and temporal) are interconnected to form a single 4D transformer block. Our complete 4D diffusion transformer architecture is composed of multiple cascaded 4D transformer blocks. Through this cascaded multi-level attention scheme, our approach significantly reduces the computational overhead while improving the training efficiency and effectively ensuring spatio-temporal consistency in the generation process.
@@
HumanArt [9] - 50k 1 $512^{2}$-$1024^{2}$ Ground Truth Half-body + Full-body
TikTok [8] 2k 600k 1 480P-720P Fitting Half-body + Full-body
TalkShow [43] 1k 300k 1 480P-720P Fitting Half-body
AIST [34] 10k 2000k 6 1080P Ground Truth Full-body
Motion-X [12] 10k 1000k 1 720P-1080P Ground Truth Full-body
DNA-rendering [4] 2k 400k 16 4K Fitting Full-body
Twindom [35] 2k 360k 180 $1024^{2}$ Fitting Full-body
THuman2.0 [44] 500 90k 180 $1024^{2}$ Ground Truth Full-body
THuman-CloSET [46] 500 90k 180 $1024^{2}$ Ground Truth Full-body
Bedlam [3] 10k 1500k 1 720P-1080P Ground Truth Full-body
@@
Human4DiT-Video 10k 2000k 1 720P-1080P Fitting Full-body

### v2

# Human4DiT: 360-Degree Human Video Generation with 4D Diffusion Transformer

@@ Abstract.
We present a novel approach for generating 360-degree high-quality, spatio-temporally coherent human videos from a single image. Our framework combines the strengths of diffusion transformers for capturing global correlations across viewpoints and time, and CNNs for accurate condition injection. The core is a hierarchical 4D transformer architecture that factorizes self-attention across views, time steps, and spatial dimensions, enabling efficient modeling of the 4D space. Precise conditioning is achieved by injecting human identity, camera parameters, and temporal signals into the respective transformers. To train this model, we collect a multi-dimensional dataset spanning images, videos, multi-view data, and limited 4D footage, along with a tailored multi-dimensional training strategy. Our approach overcomes the limitations of previous methods based on generative adversarial networks or vanilla diffusion models, which struggle with complex motions, viewpoint changes, and generalization. Through extensive experiments, we demonstrate our method’s ability to synthesize 360-degree realistic, coherent human motion videos, paving the way for advanced multimedia applications in areas such as virtual reality and animation.
journal: TOGjournal: TOGjournalyear: 2024journalvolume: 43journalnumber: 6publicationmonth: 12doi: 10.1145/3687980
@@
Figure 1. We propose Human4DiT, a novel approach to generate 360-degree high-quality, spatio-temporally coherent human videos given a reference image. With the proposed 4D diffusion transformer, our method is capable of generating monocular video, multi-view video, 3D static video, and 360-degree rotating video.
* equal contribution.
1. Introduction
@@
Human video generation is an active research topic in the field of video generation. It has widespread applications in areas such as virtual reality, animation, gaming, and movie production. Moreover, generating realistic human videos holds great significance for advancing multimedia technologies and enabling new forms of human-computer interaction.
@@
Recently, with the rapid development of diffusion models, especially latent diffusion models (Rombach et al., 2022), leveraging diffusion models for human video generation has become the mainstream approach (Wang et al., 2023; Xu et al., 2023; Zhu et al., 2024; Hu et al., 2023). To incorporate human priors as control conditions into the diffusion model, some approaches (Lee et al., 2023) have adopted skeleton-based schemes, using the skeletal connectivity graph as a control condition either through ControlNet (Zhang et al., 2023b) or by concatenating it with the input. Other methods are based on the SMPL body model template (Loper et al., 2023), injecting SMPL-derived representations such as UV maps, depth, normals (Zhu et al., 2024), or direct dense pose embeddings (Xu et al., 2023) into the diffusion model. Current human video diffusion models based on the CNNs architecture (Ronneberger et al., 2015) such as UNet from Stable Diffusion could inject control conditions into the network in a pixel-aligned manner. However, the UNet’s reliance on local convolutional operations makes it more focus on local generation, resulting in relatively poorer performance for the global aspects, especially when generating long and complex human motions. Moreover, these methods only consider the human body itself, neglecting viewpoint information from the camera perspective, especially for scenarios involving significant viewpoint changes such as 360-degree human video generation. Incorporating viewpoint control signals into the network while simultaneously maintaining coherence across different viewpoints and time poses a significant challenge.
To overcome the challenges of generating complex human motions across views and time, we propose a novel video generation network architecture that combines the strengths of CNNs and diffusion transformers. First, we propose to leverage 3D SMPL models instead of 2D skeleton maps to efficiently incorporate view information and correspondences across multiple viewpoints. Our method employs the rendering of normal human maps as robust view-dependent guidance. These rendered normal maps are subsequently processed through a CNN-based encoder, which encourages the network to capture view-dependent information and precisely inject pixel-aligned conditions.
@@
However, it remains a challenge to ensure temporal consistency when using CNNs-based architectures to generate long videos. Recently, OpenAI’s recent work on long video generation, SORA (OpenAI, 2024), adopted a diffusion transformer (Peebles and Xie, 2023) architecture and demonstrated substantially better realism and spatio-temporal coherence than CNNs-based models. Inspired by SORA, we introduce a 4D diffusion transformer for human video generation, which not only exhibits greater potential for scalability, but also demonstrates the capacity to learn complex 360-degree human video generation. By employing the unified attention mechanism across various dimensions, our 4D DiT could efficiently build spatial-temporal correspondences across different views and time, preserving spatial-temporal consistency in the generated human videos. However, directly applying a diffusion transformer to learn correlations over the views and time would be computationally prohibitive. Therefore, we propose an efficient novel 4D transformer architecture. Its core principle is to cascadingly learn the correlations across the 4D space (view, time, height, width) via self-attention. Specifically, we factorize the 4D diffusion transformer into three transformer blocks: 2D image transformer blocks, temporal transformer blocks, and view transformer blocks, each attending to different dimensions of the 4D space. These three types of blocks are interconnected to form a 4D transformer block. Multiple such 4D transformer blocks are then cascaded to construct the final 4D transformer. This efficiently captures the interrelations between body parts (height, width) across viewpoints (views) and time steps (times).
@@
To enhance the controllability of our proposed 4D diffusion transformer beyond SMPL motion, we integrate additional control signals into the respective network modules, including human identity, temporal information, and camera parameters. Human identity embeddings and latent tokens, extracted via the CLIP and a CNNs-based encoder, are incorporated into image transformers. Camera embeddings, derived from camera parameters, are integrated into the view transformer, while temporal embeddings are incorporated into the temporal transformer. Through these modules, we effectively inject various control conditions into the network, facilitating viewpoint manipulation and the generation of high-fidelity, consistent human videos.
@@
To train our proposed 4D diffusion transformer model, we also collected a large multi-dimensional dataset and devised a multi-dimensional training strategy that fully leverages all available data modalities. Our multi-dimensional dataset comprises images, videos, multi-view videos, 3D scans, as well as a limited amount of 4D scans spanning different viewpoints and time steps. During the inference stage, we propose a spatio-temporal consistent diffusion sampling strategy that enables the generation of long 360-degree videos despite limited spatio-temporal window constraints. The strategy is carried out in two stages. The first stage treats the 360-degree video as a monocular long video sequence, maximizing the temporal window to ensure long-term temporal consistency. The second stage regards the 360-degree video as a collection of multi-view video clips, using a larger viewpoint window with a smaller temporal window, encouraging consistency across viewpoints.
To summarize, our main contributions are:
@@
  • We are the first to introduce diffusion transformers to human video generation. Combined with methods that simply use CNNs-based encoder and several control modules, our method achieves high-quality 360-degree spatio-temporally consistent generation of long human videos.
@@
  • During inference, we also propose a spatio-temporal consistent diffusion sampling strategy to generate coherent 360-degree long human videos.
Figure 2. Pipeline of Human4DiT: our framework is based on 4D diffusion transformer, which adopts a cascaded structure consisting of the 2D image transformer, the view transformer, and the temporal transformer. The input contains a reference image $\mathbf{y}_{r}$, dynamic SMPL sequences $\mathbf{P}$, and camera parameters $\mathbf{c}$. Starting from a generated noisy latent representation $\mathbf{z}_{t}$, we denoise them conditioned on $\mathbf{y}_{r}$, $\mathbf{P}$, and $\mathbf{c}$ to recover the original latent frames. First, the 2D image transformer block is designed to capture spatial self-attention within latent frame tokens and pose frame tokens extracted by latent encoder $\mathbf{E}_{\mathbf{z}}$ and pose encoder $\mathbf{E}_{p}$, respectively. In addition, ID tokens $\mathbf{y}_{id}$ and image embedding $\mathbf{y}_{e}$ extracted from $\mathbf{y}_{r}$ by ID encoder $\mathbf{E}_{id}$ and CLIP are also injected to ensure identity consistency. Secondly, we use the view transformer block to learn correspondences across different viewpoints conditioned on camera embedding. Finally, we adopt a temporal transformer to capture temporal correlations with time embedding. The time embedding and camera embedding are obtained by positional encoding time $\mathbf{T}$ and camera $\mathbf{c}$, respectively.5
2. Related Work
2.1. Human Video Generation
@@
GAN-based methods (Siarohin et al., 2019b; Tian et al., 2021; Wang et al., 2021, 2020; Siarohin et al., 2018, 2019a) leverage the inherent generative capabilities of adversarial networks (Goodfellow et al., 2014; Mirza and Osindero, 2014) to spatially transform reference images according to input motion. These approaches commonly employ warping functions to generate sequential video frames, aiming to fill in missing regions and improve visually implausible areas within the generated content. While GAN-based methods have shown promising results in dynamic visual content generation, they often struggle with effectively transferring motion, especially when there are significant variations in human identity and scene dynamics between the reference image and the source video motion. This can lead to unrealistic visual artifacts and temporal inconsistencies in the synthesized videos.
@@
On the other hand, diffusion models, known for the superior generation quality and stable controllability, have been successfully integrated into human image animation. These models (Bhunia et al., 2023; Karras et al., 2023; Wang et al., 2023; Hu et al., 2023; Xu et al., 2023; Zhu et al., 2024) employ various strategies, such as texture diffusion blocks, optical flow synthesis in latent space, and motion representation using flow maps, to enhance the visual fidelity of the generated videos. Animate Anyone (Hu et al., 2023) employs a UNet-based ReferenceNet to extract features from reference images and incorporates pose information through an efficient pose guider. However, recent diffusion-based methods are still mainly relying on image diffusion model and face challenges in maintaining texture consistency and temporal stability across frames. Furthermore, these methods don’t explore view controllability.
2.2. Diffusion Model with Camera Control
@@
Camera information is conditioned in diffusion models for view control. One line of works inject camera parameters into the text-to-image diffusion models for consistent view synthesis from a single input image, which has inspired researchers to explore their potential for generating consistent multi-view images (Shi et al., 2023b; Liu et al., 2023b; kuang2024collaborative; cai2023genren). Zero-1-to-3 (Liu et al., 2023b) finetunes Stable diffusion by conditioning camera poses for zero-shot novel view synthesis. Syncdreamer (Liu et al., 2023a) employs 3D volumes and depth-wise attention to maintain consistency across views. MVDream (Shi et al., 2023b) and other methods like Wonder3D (Long et al., 2023) and Zero123++ (Shi et al., 2023a) leverage 3D self-attention to extend multi-view image generation to more general and efficient. There are other works (Yang et al., 2024; He et al., 2024) inject camera information into text-to-video (T2V) models. For example, Direct-a-Video (Yang et al., 2024) injects quantitative camera movements into temporal cross-attention blocks for view control. CameraCtrl (He et al., 2024) proposes to use an efficient representation, Plücker ray embedding, for camera conditions.
2.3. Diffusion Transformer
@@
The transformer architecture (Vaswani et al., 2017) has revolutionized the field of natural language processing, with models like GPT (Radford et al., 2018, 2019) achieving remarkable success. Recent research has demonstrated the potential of transformers in various computer vision tasks, including image classification (Touvron et al., 2021; Yuan et al., 2021), semantic segmentation (Zheng et al., 2021; Xie et al., 2021; Strudel et al., 2021).
@@
Building upon this progress, the Diffusion Transformer (DiT) (Peebles and Xie, 2023) and its variants (Bao et al., 2023; Zheng et al., 2023) have taken a step further by substituting the conventional convolutional-based U-Net backbone (Ronneberger et al., 2015) with transformers in diffusion models. This architectural shift offers enhanced scalability compared to U-Net-based models, enabling the seamless expansion of model parameters. Recently, the transformer architecture is also integrated into text-to-video models (Lu et al., 2023; OpenAI, 2024; Ma et al., 2024), improving generation performance.
3. Overview
@@
Given a reference image of a person $\mathbf{y}_{r}$, a sequence of dynamic SMPL models $\{\mathbf{P}_{i=1}^{T}=\mathbf{(\theta,\beta)}_{i=1}^{T}\}$, and camera parameters $\{\mathbf{c}_{i=1}^{V}\}$, the goal of our method is to generate a video of that person performing the corresponding motion from the specified viewpoint. The overall pipeline of our approach is illustrated in Figure 2. Since our framework is based on a latent diffusion transformer, we first generate a noisy latent representation $\mathbf{z}_{t}$. We then inject $\mathbf{y}_{r}$, $\{\mathbf{P}_{i=1}^{T}\}$, and $\{\mathbf{c}_{i=1}^{V}\}$ as control signals into the diffusion transformer, which iteratively performs denoising to ultimately generate the latent video $\mathbf{z}_{0}$ and decoded 360-degree human video $\mathbf{x}_{0}$(Sec. 4).
@@
During the inference stage, to enable the generation of 360-degree long-duration human motion videos under limited temporal and viewpoint window constraints, we propose an efficient diffusion sampling strategy. This strategy achieves improved spatial-temporal coherence by two-staged viewpoint and temporal window planning in the diffusion sampling process (Sec. 6).
Figure 3. Human4DiT dataset: In addition to the open-source dataset, we collect a multi-dimension dataset including 10k monocular videos from Internet, 5k high-quality 3D human scans and 100 animatable human models for dynamic free-view rendering.
4. Network Structure
@@
In this section, we introduce the core components of our network structure: the 4D diffusion transformer and the control condition injection modules. To efficiently establish spatial-temporal relationships, our 4D diffusion transformer adopts a cascaded structure consisting of the 2D image, the temporal, and the view transformer blocks. Additionally, our control condition injection modules comprise the reference image injection module, the SMPL injection module, the time injection module, and the viewpoint injection module. Notably, the reference image injection module and the SMPL injection module utilize CNNs-based architectures to ensure pixel-aligned injection of the control conditions.
@@
HumanArt (Ju et al., 2023) - 50k 1 $512^{2}$-$1024^{2}$ Ground Truth Half-body + Full-body
TikTok (Jafarian and Park, 2021) 2k 600k 1 480P-720P Fitting Half-body + Full-body
TalkShow (Yi et al., 2023) 1k 300k 1 480P-720P Fitting Half-body
AIST (Tsuchida et al., 2019) 10k 2000k 6 1080P Ground Truth Full-body
Motion-X (Lin et al., 2024) 10k 1000k 1 720P-1080P Ground Truth Full-body
DNA-rendering (Cheng et al., 2023) 2k 400k 16 4K Fitting Full-body
Twindom (Twindom, 2022) 2k 360k 180 $1024^{2}$ Fitting Full-body
THuman2.0 (Yu et al., 2021) 500 90k 180 $1024^{2}$ Ground Truth Full-body
THuman-CloSET (Zhang et al., 2023a) 500 90k 180 $1024^{2}$ Ground Truth Full-body
Bedlam (Black et al., 2023) 10k 1500k 1 720P-1080P Ground Truth Full-body
@@
Human4DiT-Video 10k 2000k 1 720P-1080P Fitting Half-body + Full-body